# AECBench: A Hierarchical Benchmark for Knowledge Evaluation of Large Language Models in the AEC Field

### v1

@@ Abstract
Keywords Large Language Models; Domain-specific Benchmark; Architecture, Engineering, and Construction (AEC); Knowledge Evaluation; Automated Evaluation
@@ 1 Introduction
With the ongoing intelligent transformation of the Architecture, Engineering, and Construction (AEC) industry, information technologies have played a crucial role in advancing engineering practices [1]. Among these technologies, large language models (LLMs) hold the potential to augment the knowledge-intensive tasks inherent in a building’s lifecycle, leveraging their vast knowledge repositories and cross-domain application capabilities [2, 3, 4]. To effectively harness this potential and guide the future development of LLMs for the industry, a standardized and domain-specific evaluation benchmark is essential for assessing LLMs’ proficiency in these contexts [5, 6, 7, 8].
@@
The highly specialized and safety-critical nature of the AEC domain presents a unique challenge for LLMs. As an interdisciplinary convergence of civil engineering and allied fields, the AEC domain demands extensive domain knowledge and operational capabilities spanning the entire lifecycle of a project. Figure 1 illustrates the multifaceted knowledge required in this integrative paradigm, encompassing 11 key domains, as well as numerous specialized topics radiating from these central domains. These domains are highly interrelated and collectively contribute to successful project completion. For instance, structural design is not solely governed by the principles of mechanics; it must also integrate considerations of architectural aesthetics, material properties, construction techniques, fire safety, energy efficiency, and environmental regulations, among numerous other codes and standards. Given that the AEC domain is safety-critical, this interconnectivity places stringent demands on the accuracy and reliability of LLMs. These models must be capable of assisting engineers by retrieving relevant information, facilitating cross-disciplinary decisions, and providing verifiable, evidence-based recommendations.
Recent studies have demonstrated the rapid integration of LLMs across the lifecycle of the AEC industry. In the design phase, LLMs are used to automate the generation of building energy models (BEMs) from descriptive text [9] and to enable intuitive, conversational querying of building information models (BIMs) [10]. Furthermore, LLMs are applied to automate regulatory compliance by converting textual building codes into machine-executable rules [11, 12, 13]. In structural engineering, LLMs function as controllers for the generative design of structural elements such as shear walls [14]. LLMs also facilitate structural health monitoring by analyzing sensor data [15]. During the construction phase, applications range from automating management tasks like project scheduling [16] to significantly enhancing site safety. Safety improvements include increasing the reliability of safety information retrieval by comparing retrieval-augmented generation (RAG) with fine-tuning techniques [17, 18], automatically analyzing accident reports to identify root causes [19], and developing interactive, game-based training platforms to improve worker safety knowledge [20]. These studies exemplify the potential of LLMs within the domain-specific tasks. However, they often focus on specific, narrow scenarios, leaving their broader capabilities untested. Consequently, their findings are difficult to generalize, which impedes cross-study comparability.
Advancing the robustness and reliability of LLMs in the AEC domain requires the establishment of domain-specific benchmarks. These benchmarks provide quantitative metrics to evaluate a model’s grasp of domain knowledge and its performance on relevant tasks. They play a crucial role throughout the LLM lifecycle. For practitioners, these benchmarks serve as a vital tool for model selection, offering a reliable, evidence-based means to identify the LLM best suited for their specific application needs [5, 7, 8]. For LLM developers and researchers, they provide a standardized and quantitative methodology to assess model performance, identify areas for improvement, and drive the innovation of more capable and specialized models [5, 21, 8]. Finally, for regulators, they offer the necessary empirical data to inform the creation of robust safety standards, ethical guidelines, and governance frameworks, ensuring the responsible deployment of this novel information technology in high-stakes fields [6].
Domain-specific benchmarks are being actively developed in various fields, including law, healthcare, and finance. LawBench [22] categorizes LLMs’ capabilities in law into three cognitive levels (i.e., legal knowledge memorization, understanding, and applying) and provides detailed tasks and questions for each. PromptCBLUE [23] adapts various tasks in CBLUE [24] (for example, medical named entity recognition and event extraction) into a unified language generation format to set a standard for LLMs in Chinese medical contexts. FinEVAL [25] uses finance-related multiple-choice questions to measure the effectiveness of LLMs in finance. Although benchmark frameworks developed for these fields provide valuable methodologies, they cannot be directly applied to the AEC field due to significant differences in domain characteristics. The AEC domain encompasses multidisciplinary knowledge, and the tasks require not only legal-analogous capabilities in code comprehension and logical reasoning (e.g., building code interpretation), but also finance-inspired calculation proficiency (e.g., structural load calculation), as well as unique domain-specific capabilities (e.g., structural stress analysis).
Several benchmarks have emerged to evaluate LLMs’ performance in the AEC domain. Qin et al. curated a benchmark comprising 100 fill-in-the-blank and true/false questions derived from textbook content [26], specifically from the topic "Basic Principles of Concrete Structures" within the domain of "Principles of Structural Design". Wu et al. created a larger benchmark of 875 multiple-choice questions covering a wide range of topics in the architectural domain, encompassing design, engineering, and planning knowledge [27]. These benchmarks offer valuable initial contributions to the evaluation. However, they share the limitation of a relatively narrow evaluation scope. Compared to the scope of knowledge shown in Figure 1, Qin’s benchmark covers just a part of one topic: Design of Reinforced Concrete and Steel Structures. Similarly, Wu’s benchmark is confined to the single key domain of Principles of Architectural Design. Liang et al. developed AEC-Bench, a multiple-choice evaluation consisting of over 1,000 questions sourced from official licensure examinations. This benchmark assesses abilities across twelve disciplines and five cognitive levels: knowledge memorization, understanding, reasoning, simple computation, and complex calculation [28]. Their findings reveal a progressive performance decline from lower-order cognitive tasks (memorization and understanding) to higher-order ones (reasoning and calculation). However, a key limitation of aice.AEC-Bench is its exclusive reliance on multiple-choice questions from examinations. Such questions diverge considerably from real-world AEC scenarios, where professionals are frequently required to address open-ended and ill-structured problems.
Therefore, existing evaluation benchmarks are fundamentally limited on two fronts. First, they provide an inadequate assessment of the breadth and depth of domain knowledge. Second, their misalignment with real-world scenarios means they fail to offer meaningful or trustworthy insights into a model’s practical skills. In view of these limitations, this paper presents AECBench, a comprehensive Chinese benchmark specifically designed to evaluate LLMs in real-world AEC scenarios across diverse cognitive levels. The main innovations are as follows:
@@ 2.1 Hierarchical evaluation framework
This section aims to comprehensively and effectively assess the capabilities of large language models by reflecting the unique knowledge, application scenarios, and cognitive challenges inherent to the AEC domain. Challenging the conventional approach of categorizing tasks by their disciplines or difficulties [7, 26, 27], this paper argues for a more nuanced evaluation framework for the complex AEC field. The rationale is twofold: (i) performance is dictated more by a task’s cognitive demands than its subject matter, and (ii) AEC tasks themselves are rarely cognitively monolithic. For example, a task in "Design of Reinforced Concrete and Steel Structures" necessitates memorizing code provisions, understanding load-bearing principles, reasoning about structural integrity, and creating a compliant design document. To adequately capture this complexity, this paper establishes a hierarchical framework grounded in these distinct cognitive abilities, referencing Bloom’s Taxonomy. Bloom’s Taxonomy, first introduced in 1956 and revised in 2001, is a well-established cognitive taxonomy [29, 30]. It categorizes learning objectives into six progressive levels: Remember, Understand, Apply, Analyze, Evaluate, and Create. Due to its hierarchical structure, which maps a progression from foundational knowledge to advanced problem-solving, it provides a robust theoretical basis for the proposed evaluation framework.
@@
Notably, the Knowledge Reasoning and Knowledge Calculation levels assess abilities that share conceptual similarities with those evaluated in other major benchmarks. For instance, these abilities are similar to the mathematical reasoning skills tested in prominent benchmarks like GSM8K [31] and Math-500 [32]. These conceptual similarities are also evident when compared to domain-specific benchmarks such as FinanceMath [33], MedCalc-Bench [34], and LawBench [22] for the financial, medical, and legal domains, respectively. This conceptual alignment with established benchmarks underscores the framework’s focus on critically recognized LLM competencies.
@@ 2.2 Task design
A total of 23 tasks were involved in the evaluation framework for the various cognitive levels (see Figure 2). These tasks were all derived from real-world AEC scenarios and were curated by domain engineers.
@@ 2.2.1 Knowledge Memorization tasks
This paper establishes three tasks designed to assess the memorization of key AEC knowledge, specifically targeting three typical types: (i) core law articles and code provisions, (ii) fundamental terminology, and (iii) domain-specific abbreviations.
@@ 2.2.2 Knowledge Understanding tasks
To assess the comprehension of knowledge from two critical sources, this paper establishes two types of tasks: one focused on standards and codes, and the other on established design literacy.
@@
It is noted that the code provision interpretation task is further divided into two sub-tasks, and a task (Task 2-3) that requires the model to comprehend knowledge contained within code tables is designed. Figure 3(a) illustrates an example of the question generation process from tabular data. In the figure, a multiple-choice question was crafted based on Table 6.4.5-1 of "Code for seismic design of buildings" [35], requiring the model to select the maximum axial compression ratio for a seismic wall under specific conditions.
@@ 2.2.3 Knowledge Reasoning tasks
This level includes two representative tasks designed to evaluate LLMs’ logical reasoning abilities: a deductive task that requires making design decisions based on established knowledge, and an inductive task that involves inferring a building’s type from its description.
@@ 2.2.4 Knowledge Calculation tasks
Four typical calculation tasks are involved in this level, each designed to assess quantitative abilities at a key stage of the building lifecycle: architectural and structural design, MEP design, construction, and engineering economics.
@@ 2.2.5 Knowledge Application tasks
At the Knowledge Application level, multiple representative tasks were designed for each sub-level. Five representative application tasks are considered at the Analysis sub-level: document classification, document proofreading, compliance checking, brand compliance verification, and extraction of project information. At the Evaluation sub-level, this paper primarily establishes tasks on assessing the LLM’s capabilities in acting as a rubric-based judge to score documents in architectural and structural design. At the Creation sub-level, this paper designs tasks that require LLMs to generate three typical types of documents in domain-specific workflows: conceptual design proposals, specialized reports, and tender review reports.
@@
  • Specialized Report Generation (5-3-2): Given the essential project specifications, automatically generate a structural design basis report that adheres to established codes.
@@ 2.3.1 Diverse question formats in AECBench
AECBench employs a diverse set of question formats including multiple-choice, classification, extraction, and generation. Figure 2 shows the number of questions corresponding to each individual task. Multiple-choice questions (MCQs) were used across the Knowledge Memorization, Understanding, Reasoning, and Calculation levels, as well as for the compliance checking task within the Knowledge Application level. They offer clear evaluation metrics (i.e., accuracy) and serve as a simple but effective proxy for assessing advanced capabilities of LLMs. For all tasks within the Knowledge Application level, excluding the compliance checking task, the question formats were intentionally diversified and specifically tailored to the requirements of each task. The classification format was applied to the document classification task; the extraction format was applied to the information extraction task; and a generative format was applied to both the document proofreading and brand compliance verification tasks. It must be mentioned that open-ended questions were utilized for the tasks at the Creation sub-level. Details associated with the open-ended questions are presented in Section 2.3.2.
By moving beyond a sole reliance on the multiple-choice format, the benchmark mitigates inherent limitations such as option-setting bias and the possibility of correct guessing [36]. This approach also enables a more robust evaluation of a model’s abilities in authentic AEC practice, which are difficult to assess with multiple-choice questions [37, 38].
@@ 2.3.2 Open-ended questions and automated rubric-based evaluation
A rubric-based assessment pipeline is established to facilitate the rapid and scalable evaluation of open-ended questions, where LLMs act as automated evaluation agents. This pipeline follows a two-stage workflow. Firstly, the domain experts predefine a detailed set of rubric criteria that serve as the evaluation reference. Subsequently, using the LLM-as-a-Judge approach [39], a generated document is assessed by scoring it against these expert-defined rubrics.
@@
To validate the reliability of the automated evaluation pipeline, this paper establishes parallel validation tasks in the Evaluation sub-level (Tasks 5-2-1 and 5-2-2). This involves tasks where both human experts and LLMs act as judges to score the same documents. A schematic overview of Evaluation and Creation tasks assessment methods is presented in Figure 4. In the Evaluation tasks, the model exhibiting the highest correlation with the scores of human experts is selected and subsequently employed as the judge for the generation tasks.

### v2

@@ Abstract
Keywords Large Language Models; Domain-specific Benchmark; AEC Industry; Knowledge Evaluation; Automated Evaluation
@@ 1 Introduction
The Architecture, Engineering, and Construction (AEC) industry is characterized by its inherent complexity and highly specialized, safety-critical nature. As an interdisciplinary convergence of civil engineering and allied fields, the domain demands extensive knowledge and operational capabilities spanning the entire project lifecycle [undef]. Figure 1 illustrates the multifaceted knowledge required in this integrative paradigm, encompassing 11 key domains, as well as numerous specialized topics radiating from these central domains. These domains are highly interrelated and collectively contribute to successful project completion. For instance, structural design is not solely governed by the principles of mechanics; it must also integrate considerations of architectural aesthetics, material properties, construction techniques, fire safety, energy efficiency, and environmental regulations. This deep interconnectivity places stringent demands on the accuracy and reliability of engineering practices.
Figure 1: Building lifecycle knowledge dimensions
While large language models (LLMs) offer the potential to augment these knowledge-intensive tasks by leveraging vast cross-domain repositories [undefa, undefb, undefc], deploying them in such high-stakes environments necessitates rigorous validation. Therefore, establishing a standardized and domain-specific evaluation benchmark is essential to ensure their reliability and proficiency in these contexts [undefd, undefe, undeff, undefg]. However, a comprehensive evaluation framework that fully aligns with the intricate cognitive and operational demands of the AEC industry is currently absent.
To address this critical gap, this paper presents AECBench, a comprehensive Chinese benchmark specifically designed to evaluate LLMs in real-world AEC scenarios across diverse cognitive levels. The main innovations are as follows:
@@
2 Related works
2.1 LLM integration in AEC lifecycle
Recent research demonstrates the extensive integration of LLMs across the AEC lifecycle, revealing a shift from simple queries to complex, agentic workflows. In the design phase, applications range from automating the generation of building energy models (BEMs) [undefh] to enabling conversational interactions with Building Information Models (BIMs) [undefi] and converting textual codes into executable rules for automated compliance checking [undefj, undefk]. In structural engineering, LLMs function as controllers for generative design [undefl] and analysts for structural health monitoring data [undefm]. During construction and operation, applications extend to project scheduling automation [undefn] and safety management, including accident report analysis [undefo] and retrieval-augmented safety assistance [undefp]. However, these studies currently rely on ad-hoc evaluations within specific, narrow scenarios. Consequently, there lacks a standardized benchmark to rigorously validate whether a model possesses the generalized and foundational capabilities required to support these diverse, high-stakes workflows.
2.2 Development of LLM evaluation benchmarks
Domain-specific benchmarks serve as critical infrastructure for the LLM ecosystem, offering quantitative metrics that not only validate model robustness but also guide model selection, technical optimization, and regulatory governance [undefd, undefe, undeff]. Reflecting this significance, dedicated evaluation frameworks have been established in high-stakes domains such as law, medicine, and finance [undefq, undefr, undefs]. However, the AEC industry is characterized by its unique multidisciplinary convergence, requiring not only legal-analogous code comprehension but also financial-analogous quantitative proficiency and domain-specific analysis. Consequently, the direct adaptation of frameworks from other domains is inherently insufficient.
In the specific context of the AEC domain, several initial attempts have been made to bridge this gap. Qin et al. curated a benchmark comprising 100 questions derived from textbook content [undeft], while Wu et al. created a larger dataset of 875 questions covering architectural design and planning [undefu]. Despite offering valuable initial contributions, these benchmarks suffer from fragmented coverage confined to narrow sub-domains or single topics.
To achieve more comprehensive coverage, subsequent benchmark research has gravitated towards professional licensure examinations spanning multiple disciplines within the AEC domain. Liang et al. developed aice.AEC-Bench, utilizing over 1,000 questions from licensure examinations to assess abilities across five cognitive levels [undefv]. Their findings reveal a progressive performance decline from lower-order cognitive tasks to higher-order ones. Similarly, Zhang et al. curated Civil-Eval, a collection of 426 single-choice and 91 multiple-choice items spanning eight subjects. These questions are rigorously classified into ’simple’ and ’hard’ groups [undefw]. Continuing this line of research, Zhou et al. established Civil-Bench, a larger benchmark covering 13 categories of professional engineering exam questions, including 14,823 objective questions and 269 subjective questions [undefx]. However, a key limitation persists across these benchmarks. While Civil-Bench incorporates subjective formats, it remains overwhelmingly dominated by objective problems, and aice.AEC-Bench and Civil-Eval exclusively rely on choice-based questions. Consequently, these examination-based datasets diverge considerably from real-world AEC scenarios, where professionals are frequently required to address open-ended and ill-structured problems.
@@ 3.1 Hierarchical evaluation framework
This section aims to comprehensively and effectively assess the capabilities of large language models by reflecting the unique knowledge, application scenarios, and cognitive challenges inherent to the AEC domain. Challenging the conventional approach of categorizing tasks by their disciplines or difficulties [undeff, undeft, undefu], this paper argues for a more nuanced evaluation framework for the complex AEC field. The rationale is twofold: (i) performance is dictated more by a task’s cognitive demands than its subject matter, and (ii) AEC tasks themselves are rarely cognitively monolithic. For example, a task in "Design of Reinforced Concrete and Steel Structures" necessitates memorizing code provisions, understanding load-bearing principles, reasoning about structural integrity, and creating a compliant design document. To adequately capture this complexity, this paper establishes a hierarchical framework grounded in these distinct cognitive abilities, referencing Bloom’s Taxonomy. Bloom’s Taxonomy, first introduced in 1956 and revised in 2001, is a well-established cognitive taxonomy [undefy, undefz]. It categorizes learning objectives into six progressive levels: Remember, Understand, Apply, Analyze, Evaluate, and Create. Due to its hierarchical structure, which maps a progression from foundational knowledge to advanced problem-solving, it provides a robust theoretical basis for the proposed evaluation framework.
@@
Notably, the Knowledge Reasoning and Knowledge Calculation levels assess abilities that share conceptual similarities with those evaluated in other major benchmarks. For instance, these abilities are similar to the mathematical reasoning skills tested in prominent benchmarks like GSM8K [undefaa] and Math-500 [undefab]. These conceptual similarities are also evident when compared to domain-specific benchmarks such as FinanceMath [undefac], MedCalc-Bench [undefad], and LawBench [undefq] for the financial, medical, and legal domains, respectively. This conceptual alignment with established benchmarks underscores the framework’s focus on critically recognized LLM competencies.
@@ 3.2 Task design
To ensure the comprehensive representativeness of the benchmark, the 23 tasks were selected to reflect both the breadth of the AEC project lifecycle and the depth of cognitive skills required in the field. The tasks capture breadth by systematically covering core professional activities from conceptual design and regulatory management to engineering and construction. Simultaneously, they capture depth by mapping these activities to the five cognitive levels (see Figure 3)—ranging from foundational knowledge to complex application. These tasks were all derived from real-world AEC scenarios and were curated by domain engineers to ensure practical relevance.
@@ 3.2.1 Knowledge Memorization tasks
This paper establishes three tasks designed to assess the memorization of key AEC knowledge, specifically targeting three typical types: (i) core law articles and code provisions, (ii) fundamental terminology, and (iii) domain-specific abbreviations. These three types were selected for their universal representativeness: Core law articles and code provisions constitute the industry’s mandatory regulatory and safety bedrock. Fundamental terminology provides the essential vocabulary for all unambiguous technical communication between architects, engineers, and contractors. Domain-specific abbreviations are the high-frequency, practical shorthand used universally across all project documentation, from technical drawings to reports.
@@ 3.2.2 Knowledge Understanding tasks
To assess the comprehension of knowledge from two critical sources, this paper establishes two types of tasks: one focused on standards and codes, and the other on established design literacy. The comprehension of standards and codes tests the core professional skill of interpreting specific requirements, both from complex text and from the dense tabular data ubiquitous in engineering documents. Established design literacy represents the complementary, fundamental body of engineering principles and concepts required for practical application.
@@
It is noted that the code provision interpretation task is further divided into two sub-tasks, and a task (Task 2-3) that requires the model to comprehend knowledge contained within code tables is designed. Figure 4(a) illustrates an example of the question generation process from tabular data. In the figure, a multiple-choice question was crafted based on Table 6.4.5-1 of "Code for seismic design of buildings" [undefae], requiring the model to select the maximum axial compression ratio for a seismic wall under specific conditions.
@@ 3.2.3 Knowledge Reasoning tasks
To evaluate LLMs’ abilities of reasoning with knowledge to make sound, compliant judgments, this level includes two representative tasks designed to evaluate LLMs’ logical reasoning abilities: a deductive task that requires making design decisions based on established knowledge, and an inductive task that involves inferring a building’s type from its description. Deductive reasoning (Design Decision Formulation) is the critical process of applying complex rules and standards to specific project scenarios. This skill is paramount as it directly impacts engineering safety and code compliance. Concurrently, inductive reasoning (Building Type Inference) assesses the vital diagnostic skill of making preliminary judgments about a project. This capacity is a fundamental prerequisite for countless downstream tasks, as professionals must first correctly identify a project’s type (e.g., residential, hospital, industrial) before they can apply the correct set of standards or make sound design decisions.
@@ 3.2.4 Knowledge Calculation tasks
To assess the quantitative ability to execute precise, verifiable calculations, the four selected tasks were designed to be comprehensively representative, capturing how this essential skill manifests ubiquitously across the entire building lifecycle. This includes the core calculations required in architectural and MEP design, the practical mathematics of the construction phase, and the fundamental financial analysis of engineering economics. These tasks collectively assess the critical numerical literacy required in the industry.
@@ 3.2.5 Knowledge Application tasks
At the Knowledge Application level, multiple representative tasks were designed for each sub-level. Five representative application tasks are considered at the Analysis sub-level: Document Classification, Document Proofreading, Compliance Checking, Brand Compliance Verification, and Information Extraction. These tasks assess the essential, high-frequency skills that define daily practice. This includes not just managing information flow by classifying and categorizing documents or extracting entities from tenders, but the critical, high-stakes processes of performing calculations and reasoning to check code compliance and detecting brand-related deviations in bid submissions. At the Evaluation sub-level, the tasks assess the fundamental competency of acting as a rubric-based judge to score professional documents, a vital simulation of the industry’s critical peer-review and quality assurance loops. At the Creation sub-level, this paper designs tasks that require LLMs to generate high-value, structured documents that propel a project, such as the conceptual design proposals that initiate the planning phase, the specialized reports that define the detailed design, and the tender review reports from the procurement phase that summarize discrepancies.
@@
  • Specialized Report Generation (5-3-2): Given the essential project brief, automatically generate a structural design basis report that adheres to established codes.
@@ 3.3.1 Diverse question formats in AECBench
AECBench employs a diverse set of question formats including multiple-choice, classification, extraction, and generation. Figure 3 presents the distribution of questions across individual tasks, which are hierarchically organized by cognitive level and color-coded to denote these distinct formats. Multiple-choice questions (MCQs) were used across the Knowledge Memorization, Understanding, Reasoning, and Calculation levels, as well as for the compliance checking task within the Knowledge Application level. They offer clear evaluation metrics (i.e., accuracy) and serve as a simple but effective proxy for assessing advanced capabilities of LLMs. For all tasks within the Knowledge Application level, excluding the compliance checking task, the question formats were intentionally diversified and specifically tailored to the requirements of each task. The classification format was applied to the document classification task; the extraction format was applied to the information extraction task; and a generative format was applied to both the document proofreading and brand compliance verification tasks. It must be mentioned that open-ended questions were utilized for the tasks at the Creation sub-level. Details associated with the open-ended questions are presented in Section 3.3.2.
By moving beyond a sole reliance on the multiple-choice format, the benchmark mitigates inherent limitations such as option-setting bias and the possibility of correct guessing [undefaf]. This approach also enables a more robust evaluation of a model’s abilities in authentic AEC practice, which are difficult to assess with multiple-choice questions [undefag, undefah].
@@ 3.3.2 Open-ended questions and automated rubric-based evaluation
A rubric-based assessment pipeline is established to facilitate the rapid and scalable evaluation of open-ended questions, where LLMs act as automated evaluation agents. Compared to the reference-free evaluation, this rubric-based pipeline anchors the LLM’s judgment to objective, expert-defined criteria. This methodology has been suggested to ensure the evaluation is a trustworthy indicator of domain expert opinion, enabling the model-based evaluator to achieve a level of reliability comparable to inter-annotator agreement among human experts [undefah, undefai]. This pipeline follows a two-stage workflow. Firstly, the domain experts predefine a detailed set of rubric criteria that serve as the evaluation reference. Subsequently, using the LLM-as-a-Judge approach [undefai], a generated document is assessed by scoring it against these expert-defined rubrics.
@@
To validate the reliability of the automated evaluation pipeline, this paper establishes parallel validation tasks in the Evaluation sub-level (Tasks 5-2-1 and 5-2-2). This involves tasks where both human experts and LLMs act as judges to score the same documents. In the Evaluation tasks, the model exhibiting the highest correlation with the scores of human experts is selected and subsequently employed as the judge for the generation tasks. A schematic overview of the assessment framework for Evaluation and Creation tasks is presented in Figure 5.